Develop and present a valuation model for corporate debt with a face value of $100 million dollars. The model should use hypothetical assumptions for the coupon rate and other characteristics as well as a hypothetical market interest rate. You must also select a maturity for the bonds and the frequency of the coupon payments. The market rate should be justifiable/reasonable given current market conditions.
Answer #1

Developing a bond for corporate debt
Par Value =100 million dollars
Assumptions
Number of Frequency =2 Coupons per year
Number of Coupons per year =2(Semi annual coupon)
Coupon rate =8%
Hence Semi annual Coupon =8%/2 =4%
Coupon =4%*1000 =40
Let YTM =5%
Semi annual YTM =5%/2 =2.5%
Number of Years of Bond =5
Number of Periods =5*2 =10
Valuation model using for bond
(Valuation model for bond )Price of Bonds =PV of Coupons +PV of Par Value =Coupons*((1-(1+YTM/2)^-n)/(YTM/2) +Par Value/(1+YTM)^n =40*((1-(1+2.5%)^-10)/2.5%)+1000/(1+2.5%)^10 =1131.281